Two flat mirrors are perpendicular to each other. An incoming beam of light makes an angle of 15° with the first mirror as shown below. What angle will the outgoing beam make with the surface of the second mirror?


1
Expert's answer
2021-07-07T08:39:29-0400

"θ_1=θ_2 = 90 -15 = 75° \\

θ_3= 180 -90 -75 = 15° \\

θ_3=θ_4 \\

θ_5 = 90 -θ_4 = 90 -15 = 75°"

Answer: 75°